The triangle inequality states that for any triangle, the sum of the lengths of any two sides must be greater than or equal to the length of the remaining side.
Calculation:
Given: a, b, c are the sides of a triangle ABC
Since, a, b, c are positive numbers.
⇒ a1/p, b1/p, c1/p are positive numbers. ...(∵ p > 1)
We know that sum of any two sides must be greater than third side.
⇒ a1/p + b1/p > c1/p
⇒ a1/p + b1/p - c1/p > 0
∴ Option 2 is correct.
For the next two (02) items that follow:
Consider the function f(x) = -2x3 – 9x2 – 12x + 1
Question 6
The function f(x) is an increasing in the interval
The intercept form of the plane is given by ⇔ ax+by+cz=1
Where ‘a’ is the x- intercept, ‘b’ is the y- intercept and ‘c’ is the z- intercept.
Calculation:
Given:
Equation of plane 5x + 2y + z – 13 = 0
⇒ 5x + 2y + z = 13
⇒135x;+;2y;+;z;=1
⇒513x+213y+113z=1
∴ Lengths of intercepts are 513,213,;13 unit
Question 9
The number of ways in which 3-holiday tickets can be given to 20 employees of an organization if each employee is eligible for any one or more of the tickets, is
Whenever the situations like finding the possible arrangements for m slots given to n people and there is no such restriction of giving the particular number of slots to any of their person, then a total number of possible arrangements is given by = m × m × m.....(n times) = mn .
Calculation:
Since, In the given question, Each ticket can be given to 20 employees and there are 3 tickets in total.
Thus, each of the 3 tickets has 20 different options.
i.e. There are the following possible ways:
20 × 20 × 20 = 203 = 8000.
Question 10
Seven unbiased coins are tossed 128 times. In how many throws would you find at least three heads?
Let us consider the standard form of a quadratic equation, ax2 + bx + c =0
Where a, b and c are constants with a ≠ 0
Let α and β be the two roots of the above quadratic equation.
The sum of the roots of a quadratic equation are:
α+β=−ab=−coefficient;of;x2coefficient;of;x
The product of the roots is given by:
αβ=ac=coefficient;of;x2constant;term
Calculation:
We have, ax2 + bx + c =0
Let α and β be the two roots of the above quadratic equation.
The sum of the roots of a quadratic equation are: α+β=−ab
The product of the roots is given by: αβ=ac
Given: Sum of the roots of the equation is equal to the sum of their squares.
∴ α + β = α2 + β2
⇒ α + β = (α + β) 2 - 2αβ
⇒−ab=;(−ab)2−2;×ac
⇒−ab=a2b2−a2c
⇒−ab=a2b2−2ac
⇒ -ab = b2 -2ac
∴ b2 + ab = 2ac
Question 14
Consider the following statements in respect of the determinant \(\left| {\begin{array}{*{20}{c}} {{{\cos }^2}\frac{\alpha }{2}}&{{{\sin }^2}\frac{\alpha }{2}}\ {{{\sin }^2}\frac{\beta }{2}}&{{{\cos }^2}\frac{\beta }{2}} \end{array}} \right|\)
If \({\rm{A}} = \left[ {\begin{array}{*{20}{c}} {{{\rm{a}}{11}}}&{{{\rm{a}}{12}}}\ {{{\rm{a}}{21}}}&{{{\rm{a}}{22}}} \end{array}} \right]\) then determinant of A is given by: |A| = (a11 × a22) - (a12 - a21).
Where byx and bxy are regression coefficients or the slopes of the equation y on x and x on y are denoted as byx and bxy
Note: If bxy and byx are negative then the coefficient of correlation would be negative and if If bxy and byx are positive then the coefficient of correlation would be positive.
An insurance company insured 2000 scooter drivers, 4000 car drivers and 6000 truck drivers. The probabilities of an accident involving a scooter driver car driver and a truck driver are 0.01, 0.03 and 0.15 respectively. One of the insured persons meets with an accident. The probability that the person is a scooter driver is
A problem in statistics is given to three students A, B and C whose chances of solving it independently are 21,31 and 41 respectively. The probability that the problem will be solved is
The complement of an event is the subset of outcomes in the sample space that are not in the event. A complement is itself an event.
The probability of the complement of an event is one minus the probability of the event.
P (A’) or P (Ac) or P(Aˉ) = 1 - P (A)
Where,
P (A) be probability of Event A and P (A’) or P (Ac) or P(Aˉ) be probability of the complement of Event A
Calculation:
Given:
Probability of solving problem by A = P (A) =1/2
Probability of solving problem by B = P (B) =1/3
Probability of solving problem by C = P (C) =1/4
P (problem will not solved by A) = 1 - (1/2) = 1/2
P (problem will not solved by B) = 1 - (1/3) = 2/3
P (problem will not solved by C) = 1 - (1/4) = 3/4
P (Problem will be solved) = 1 - P (problem will not solved by A, B and C)
=1−(21×32×43)=1−41=43
Question 22
A bag contains 4 white and 2 black balls and another bag contains 3 white and 5 black balls. If one ball is drawn from each bag, then the probability of one ball being white and the other ball being black is
Equation of a line perpendicular to a given line ax + by + c = 0 is bx - ay + λ = 0, where λ is a constant.
The intercept form of the line ⇔ ;ax+by=1 , Where a is the x- intercept and b is the y- intercept.
Calculation:
Given: Area of triangle is 5 square units
Given equation of line is 5x – y = 0
Equation of a line perpendicular to a given line is
⇒ x + 5y = λ …. (1)
⇒λx+;λ5y=1⇒λx+;(5λ)y=1
Area of triangle = 5
⇒21×λ×5λ=5
⇒ λ2 = 50
∴ λ = ± 5√2
Put the value of λ in equation 1st
So, x + 5y = ± 5√2
⇒ x + 5y ± 5√2 = 1
Question 28
Consider any point P on the ellipse 25x2+9y2=1 in the first quadrant. Let r and s represent its distance from (4, 0) and (-4, 0) respectively, then (r + s) is equal to
Ellipse: The locus of a point which moves such that its distance from a fixed point and a fixed straight line is always less than one (Eccentricity = e < 1) ⇔ a2;x2+b2y2=1 (a > b)
The sum of the focal distance of any point on an ellipse is constant and equal to the length of the major axis of the ellipse. ⇔ PS + PS' = 2a = Major axis
Focus: (± ae, 0)
Eccentricity: e=;1−;a2b2
Calculation:
We have 25x2+9y2=1
Here, a2 = 25, b2 = 9 ⇒ a = 5 and b = 3
Eccentricity:
e=;1−;a2b2=;1−;259=;2516=;54
Focus = (± ae, 0) = (±4, 0)
PS + PS' = 2a = Major axis
⇒ r + s = 2 × 5 = 10 unit
Question 29
A straight line x = y + 2 touches the circle 4(x2 + y2) = r2. The value of r is
Hyperbola: The locus of a point which moves such that its distance from a fixed point is greater than its distance from a fixed straight line. (Eccentricity = e > 1)
Order: The order of a differential equation is the order of the highest derivative appearing in it.
Degree: The degree of a differential equation is the power of the highest derivative occurring in it, after the Equation has been expressed in a form free from radicals as far as the derivatives are concerned.
Calculation:
Given:
dxdy−x=(y−xdxdy)−4
⇒(dxdy−x)=(y−xdxdy)41 (∵ x-1 = -1/x)
⇒(dxdy−x)(y−xdxdy)4=1
⇒dxdy×(y−xdxdy)4−x×(y−xdxdy)4=1
⇒dxdy×[(xdxdy)4−…]−x×(y−xdxdy)4=1
⇒x4(dxdy)5….=1
We know that, the degree of a differential equation is the power of the highest derivative
Equation of parabola having vertex at the origin and focus at (a, 0), where a > 0, is y2 = 4ax
Order: The order of a differential equation is the order of the highest derivative appearing in it.
Degree: The degree of a differential equation is the power of the highest derivative occurring in it, after the Equation has been expressed in a form free from radicals as far as the derivatives are concerned.
Calculation:
Equation of parabola having vertex at the origin and focus at (a, 0), where a > 0, is y2 = 4ax
y2 = 4ax ---(1)
Differentiating both sides, we get
⇒2ydxdy=4a
On putting the value of (4a) in eq. (1) we get
y2=2yxdxdy⇒y=2xdxdy
⇒2x(dxdy)−y=0
Hence, the order and degree of differential equation are 1 and 1, respectively
i.e. Left hand limit = right hand limit = value at the function.
⇒x→2π−limf(x)=x→2π+limf(x)=f(2π)
⇒ (A + B) = 0 = 0
⇒ A + B = 0..........(equation 1)
Now,
f(x) is also continuous at x = - 2π
i.e. Left hand limit = right hand limit = value at the function.
⇒x→2−π−limf(x)=x→2−π+limf(x)=f(−2π)
⇒ 2 = ( - A + B) = 2
⇒ - A + B = 2..........(equation 2)
On solving equation 1 and 2, we get
A = -1 and B = 1
Thus, the value of B = 1.
Question 44
Let X be the set of all persons living in Delhi. The person's a and b in X are said to be related if the difference in their ages is at most 5 years. The relation is?
2. Symmetric: If any one element is related to any other element, then the second element is related to the first.
R is symmetric if for all a, b ∈ X, if aRb, then bRa.
3. Transitive: If any one element is related to a second and that second element is related to a third, then the first element is related to the third.
R is transitive if for all a, b, c ∈ X, if aRb and bRc, then aRc.
R is an equivalence relation if A is nonempty and R is reflexive, symmetric and transitive.
Calculation:
Given: X = {Set of persons living in Delhi}
R = {(a, b) : ∣a - b∣ ≤ 5}
For reflexive,
Let (a, a) ∈ R
(a, a) = ∣a - a∣ = 0,
So, 0 ≤ 5 ⇒ (a, a) ≤ 5
Hence, R is reflexive.
For symmetric,
Let (a, b) ∈ R
(a, b) = ∣a - b∣ ≤ 5
⇒ ∣b - a∣ ≤5
⇒ (b, a) ∈ R
∴ (a, b) ∈ R ⇒ (b, a) ∈ R
Hence, R is symmetric.
For transitive,
Let (a, b) ∈ R and (b, c) ∈ R
(a, b) = ∣a - b∣ ≤ 5, and (b, c) = ∣b - c∣ ≤ 5
Now,
∣a - c∣ ≤ ∣a - b∣ + ∣b - c∣ ≤ 5 + 5 ≤ 10
⇒ ∣a - c∣ ≤ 10
∴ (a, c) ∉ R
Hence, R is not transitive.
∴ R is reflexive and symmetric but not transitive.
If the value of the determinant \(\left[ {\begin{array}{*{20}{c}} {\rm{a}}&1&1\ 1&{\rm{b}}&1\ 1&1&{\rm{c}} \end{array}} \right]\) is positive, where a ≠ b ≠ c, then the value of abc is
Let us consider sequence a1, a2, a3 …. an is an G.P.
Common ratio = r = \(\frac{{{{\rm{a}}_2}}}{{{{\rm{a}}_1}}} = \frac{{{{\rm{a}}3}}}{{{{\rm{a}}2}}} = \ldots = \frac{{{{\rm{a}}{\rm{n}}}}}{{{{\rm{a}}{{\rm{n}} - 1}}}}\)
nth term of the G.P. is an = arn-1
Sum of n terms = s = r−;1a;(rn−1); where r >1
Sum of n terms = s = 1−;ra;(1−;rn); where r <1
Sum of infinite GP = s∞=1;−;ra; |r| < 1
Calculation:
Let;S=621×621×683×641×…
This can be written as:
S=621×642×683×6164×…
S=621+42+83+164+…
Let x=21+42+83+164+… ---(1)
Multiply by 1/2 on both sides, we get:
⇒2x=41+82+163+324+… ---(2)
Subtract equation 2 from equation 1, we get
x−2x=(21+42+83+164+…)−;(41+82+163+324+…)
⇒2x=21+(42−41)+(83−82)+(164−163)+…
⇒2x=21+41+81+161+321+…
The above series is clearly a Geometric Progression with the first term = 1/2 and common ratio = 1/2
∴2x=(1−21)21=2121=1
⇒ x = 2 × 1 = 2
Put the value of x in the original sum:
∴ S = 62 = 36
Question 51
If the roots of the equation x2 - nx + m = 0, Differ by 1, then
Let us consider the standard form of a quadratic equation, ax2 + bx + c =0
Let α and β be the two roots of the above quadratic equation.
The sum of the roots is given by: α+β=−ab=−coefficient;of;x2coefficient;of;x
The product of the roots is given by: αβ=ac=coefficient;of;x2constant;term
Calculation:
We have, x2 - nx + m = 0
Let α and β be the two roots of the above quadratic equation.
Sum of the roots = α + β = n
Product of the roots = αβ = m
Given: Roots are differ by 1,
∴ α - β = 1
We know that,
(α + β)2 = (α - β)2 + 4 αβ
⇒ n2 = 12 + 4m
∴ n2 - 4m - 1 = 0
Question 52
If different words are formed, with all the letters of the word ‘AGAIN’ and are arranged alphabetically among themselves as in a dictionary, the word at the 50th place will be
Suppose a word of n letters has n1 of one kind of letter, n2 of a second kind, n3 of a third kind, and so on with n = n1 + n2 + n3 + ... + nk then the number of distinguishable permutations of the n letters is = \(\frac{{{\rm{n}}!}}{{{{\rm{n}}_1}! \times {{\rm{n}}_2}! \times {{\rm{n}}3}!...{{\rm{n}}{\rm{k}}}!}}\)
Calculations:
We have, word ‘AGAIN’
Arrange alphabetically: AAGIN
The letter starts from A,
A
A
G
I
N
AAGIN = 4! = 24
The letter starts from G,
G
A
A
I
N
GAAIN = (4!)/2 = 12
The letters starts from I,
I
A
A
G
N
IAAGN = (4!)/2 = 12
Number of total letters = 24 + 12 + 12 = 48
49th word is NAAGI
50th word is NAAIG.
Question 53
The number of ways in which a cricket team of 11 players can be chosen out of a batch of 15 players so that the captain of the team is always included, is
Suppose a word of n letters has n1 of one kind of letter, n2 of a second kind, n3 of a third kind, and so on with n = n1 + n2 + n3 + ... + nk then the number of distinguishable permutations of the n letters is = \(\frac{{{\rm{n}}!}}{{{{\rm{n}}_1}! \times {{\rm{n}}_2}! \times {{\rm{n}}3}!...{{\rm{n}}{\rm{k}}}!}}\)
The number of ways of selecting r objects from n different objects subject to certain conditions like:
k particular objects are always included = n - kCr - k
k particular objects are never included = n - kCr
Calculation:
We know that The number of ways of selecting r objects from n different objects subject to a certain condition is:
k particular objects are always included = n - kCr - k
∴ Total number of selection of 11 players out of 15 players in which the captain is included = 14C10 = 1001
Question 54
In the expansion of (x+3x21)10 the value of constant term (independent of x) is
Two poles are 10 m and 20 m high. The line joining their tops makes an angle of 15° with the horizontal. The distance between the poles is approximately equal to
If a, b, c are real numbers, then the value of the determinant \(\left| {\begin{array}{*{20}{c}} {1 - {\rm{a}}}&{{\rm{a}} - {\rm{b}} - {\rm{c}}}&{{\rm{b}} + {\rm{c}}}\ {1 - {\rm{b}}}&{{\rm{b}} - {\rm{c}} - {\rm{a}}}&{{\rm{c}} + {\rm{a}}}\ {1 - {\rm{c}}}&{{\rm{c}} - {\rm{a}} - {\rm{b}}}&{{\rm{a}} + {\rm{b}}} \end{array}} \right|\) is
Binomial distribution The binomial distribution formula is:P (X = r) = \({{\bf{n}}{{{\bf{C}}{\bf{r}}}}}{{\bf{p}}^{\bf{r}}}{{\bf{q}}^{{\bf{n}} - {\bf{r}}}}\)
P(X) = probability of a success on an individual trial
Binomial distribution The binomial distribution formula is:P (X = r) = \({{\bf{n}}{{{\bf{C}}{\bf{r}}}}}{{\bf{p}}^{\bf{r}}}{{\bf{q}}^{{\bf{n}} - {\bf{r}}}}\)
P(X) = probability of success on an individual trial
The geometric mean of the observations x1, x2, x3, … xn is G1. The geometric mean of the observations y1, y2, y3,… yn is G2. The geometric mean of observations \(\frac{{{{\rm{x}}_1}}}{{{{\rm{y}}_1}}},\frac{{{{\rm{x}}_2}}}{{{{\rm{y}}_2}}},\frac{{{{\rm{x}}3}}}{{{{\rm{y}}3}}}, \ldots \frac{{{{\rm{x}}{\rm{n}}}}}{{{{\rm{y}}{\rm{n}}}}}\) is
The geometric mean of the observations x1, x2, x3, … xn is G1
∴ G1 = (x1 x2 x3 … xn)(1/n) …. (1)
Similarly,
G2 = (y1 y2 y3 … yn)(1/n) …. (2)
Now, the geometric mean of observations \(\frac{{{{\rm{x}}_1}}}{{{{\rm{y}}_1}}},\frac{{{{\rm{x}}_2}}}{{{{\rm{y}}_2}}},\frac{{{{\rm{x}}3}}}{{{{\rm{y}}3}}}, \ldots \frac{{{{\rm{x}}{\rm{n}}}}}{{{{\rm{y}}{\rm{n}}}}}\)
An unbiased coin is tossed until the first head appears or until four tosses are completed, whichever happens earlier. Which of the following statements is/are correct?
The probability that no head is observed is 161.
The probability that the experiment ends with three tosses is 81.
Select the correct answer using the code given below:
So, in the composite functions a an odd number of g(x) will result in reciprocal of x, whereas an even number of g(x) results in x itself. The number of f does not cause any difference.
i.e f(x) gives the function itself and g(x) gives reciprocal
Question 84
If A is an orthogonal matrix of order 3 and \({\rm{B}} = \left[ {\begin{array}{*{20}{c}} 1&2&3\ { - 3}&0&2\ 2&5&0 \end{array}} \right]\), then which of the following is/are correct?
|AB| = ± 47
AB = BA
Select the correct answer using the code given below:
In any polynomial of n sides, the number of diagonals are as follows:
For making a diagonal each vertex can be connected to n – 3 vertices, except for both adjacent vertices and itself. The same process repeated for all n vertices gives a total of n(n - 3)
Where each diagonal is counted twice for each of its vertex.
Number;of;diagonals=2n(n−3)
Calculation:
Number of diagonals:
2n(n−3)=44
⇒ n2 - 3n = 88
⇒ n2 - 3n – 88 = 0
⇒ n(n - 11) + 8(n - 11) = 0
⇒ (n + 8)(n - 11) = 0
Thus, possible values of n are n = -8 and n = 11.
Since number of sides of a polygon cannot be negative, n = 11
Question 87
If p, q, r are in one geometric progression and a, b, c are in another geometric progression, then ap, bq, cr are in
General term: General term in the expansion of (x + y)n is given by: T(r+1)=;nCr×xn−r×yr
Middle terms: The middle terms is the expansion of (x + y)n depends upon the value of n.
If n is even, then total number of terms in the expansion of (x + y)n is n + 1. So there is only one middle term i.e. (2n+1)th term is the middle term.
If n is odd, then total number of terms in the expansion of (x + y)n is n + 1. So there are two middle terms i.e. (2n+1)th and (2n+3)th are two middle terms.
3. ncr+ncr+1;=;n+1cr+1
4. ncr=rnn−1cr−1
Calculation:
We have to find average of the coefficients of the two middle terms in the expansion of (1 + x) 2n+1
Since, 2n + 1 is odd.
Hence (22n+1+1);and;(22n+1+3) are two middle
⇒ (n + 1)th and (n + 2)th term are two middle terms.
Now, coefficients of the (n + 1)th term and (n + 2)th term are 2n+1cn and 2n+1cn+1
Average of the coefficients of the two middle terms = 22n+1cn+2n+1cn+1
We know that ncr+ncr+1=n+1cr+1
So, 2n+1cn+2n+1cn+1=2n+2cn+1
Average of the coefficients of the two middle terms = 22n+1cn+2n+1cn+1=22n+2cn+1
We know that ncr=rnn−1cr−1
So, 2n+2cn+1=n+12n+22n+1cn=n+12(n+1)2n+1cn=2×2n+1cn
Average of the coefficients of the two middle terms = 22n+2cn+1=22×2n+1cn=2n+1cn
∴ Option 2 is correct.
Question 91
The sum of the coefficients of all the terms in the expansion is
We have to calculate sum of the coefficients of all the terms in the expansion of (1 + x) 2n+1
⇒ Sum of the coefficients = 2n+1c0+2n+1c1+2n+1c2+…+2n+1c2n+1
⇒ Sum of the coefficients = 2(2n + 1) = 22n × 2 = 2 × 4n
Alternate solution:
To find the sum of coefficient of all terms, put x = 1 in the given expression (1 + x) 2n+1, we get
Sum of the coefficients = 2(2n + 1) = 22n × 2 = 2 × 4n
Question 92
Two straight lines passing through the point A(3, 2) cut the line 2y = x + 3 and x-axis perpendicularly at P and Q respectively. The equation of the line PQ is
Let line l = xi + yj +zk is perpendicular to lines l1 = x1i + y1j + z1k and l2 = x2i + y2j + z2k then l is the cross product of other two lines i.e, l = l1 × l2
L-Hospital Rule: Let f(x) and g(x) be two functions
Suppose that we have one of the following cases,
I. x→alimg(x)f(x)=;00
II. x→alimg(x)f(x)=;∞∞
Then we can apply L-Hospital Rule ⇔ x→alimg(x)f(x)=;x→alimg′(x)f′(x)
Note: We have to differentiate both the numerator and denominator with respect to x unless and until x→alimg(x)f(x)=l=00 where l is a finite value.
Complex number z = x + iy, on graph represent x = x-coordinate y = y-coordinate where Y-axis is imaginary axis and X-axis is real axis.
Calculation:
Let, z = x + iy ⇒ z̅ = x - iy
iz̅ - iz = 5 will be
⇒ i (x - iy) - i (x + iy) = 5
⇒ xi - i2y - ix - i2y = 5
⇒ 2y = 5 (∵ i2 = -1)
⇒ y = 5/2
x - coordinate of z2 will be the same as of the z1 as the point z2 is the reflection of the z1 about the line y = 5/2
∴ x-coordinate of z2 = 1
Distance |Cz1| = (√ [(1 - 1)2 + (5/2 - 1)2)
⇒ Distance |Cz1| = 3/2
Now, y-coordinate of z2 = 5/2 + 3/2 = 8/2
⇒y-coordinate of z2 = 4
∴ z2 = 1 + 4i
Hence, option (1) is correct.
Question 110
A vertical tower standing on a levelled field is mounted with a vertical flag staff of length 3 m. From a point on the field, the angles of elevation of the bottom and tip of the flag staff are 30° and 45° respectively. Which one of the following gives the best approximation to the height of the tower?